OpenStudy (anonymous):

Find the Surface area and volume of a cone with the diameter of 6 and the height of 5

OpenStudy (anonymous):

SA = πr2 + πrl SA = Surface Area, use the formula to find the Surface Area. Not sure, about the volume of a cone, sorry.

OpenStudy (anonymous):

Okay, Another question how would I find the slant thingy

OpenStudy (anonymous):

the slant height?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Oh shoot I sounded stupid. Nvm I got it. Thank you

OpenStudy (anonymous):

radius^2+height^2=c^2 (act like, the radius was 10, height with 24) 10^2+24^2=c^2 (10^2 = 100 ; 24^2 = 576) 100+576=c^2 676 = c^2 c=26 <--- slant height.

OpenStudy (anonymous):

and okay, not sure If I had it right, but welcome. ;P